# 2]organization=Shanghai Blockchain Engineering Research Center, city=Shanghai, postcode=200433, country=China

### v1

@@ 4 UIFV Framework Overview
We have developed a comprehensive framework, the Unified InverNet Framework in VFL (UIFV), which enables attackers with varying capabilities to train an InverNet and use it to reconstruct private data.
In the VFL framework, we consider the bottom model $f_{k}$ of participant $\mathbb{P}_{k}$ as a feature extractor. The bottom model $f_{k}$ generates the intermediate features $H^{k}$ that are fed into the top model $f_{\text{top}}$. Considering that the intermediate layers of neural networks retain rich semantic information of input data, the proposed UIFV framework focuses on training an InverNet $g$ that establishes the relationship $g(f_{k}(x^{k}))=x^{k}$, which can then be used to upsample and reconstruct the private data $x^{k}$. The objective function for training the InverNet $g$ can be formulated as:
$\arg\min_{\theta_{g}}\left\|g(f_{k}(x^{k}))-x^{k}\right\|^{2},$ (2)
which indicates that the private data $x^{k}$ and the bottom model $f_{k}$ are required for training the InverNet. However, such information cannot be directly obtained in realistic VFL scenarios. Therefore, Data Preparation and Model Preparation are two key functional modules in the proposed UIFV framework that respectively prepare the model and data information needed for training the InverNet.
@@
Similarly, different methods can be employed by the Model Preparation module based on the attacker’s capabilities. If the attacker can make inference queries to the target party, which is reasonable in the VFL architecture, then the model output $f_{k}(x^{k})$ can be obtained from the queries. If the attacker cannot make query requests to the bottom model, then the attacker may train a shadow model $\tilde{f}_{k}$ to replace the model $f_{k}$ for training the InverNet.
@@
As depicted in Figure 2, the UIFV framework encompasses three distinct stages. The Data Preparation and Model Preparation functions are performed in the initial stage to prepare the pertinent data and model needed for training the InverNet, which is then conducted in the second stage. Then, in the final stage, the pretrained InverNet is employed to reconstruct the private data used for model training by the target party.
@@ 5.1 Query Attack (QA)
In the Query Attack scenario, the attacker has no access to the target party’s model parameters and model gradient information during the training process, which is consistent with the black-box attack setting Jiang et al. (2022); Liu et al. (2021). On the other hand, the attacker can freely initiate query requests to the target party and possesses the auxiliary dataset $\tilde{x}^{\text{target}}$, that is, i.i.d. with the target party’s private data.
$g=\arg\min_{\theta_{g}}\frac{1}{m}\sum_{i=1}^{m}\left\|g(\tilde{H}^{\text{target}}_{i})-\tilde{x}^{\text{target}}_{i}\right\|^{2}$ (3)
In this attack scenario, since the attacker owns an auxiliary data set and is able to query the target model, the UIFV framework requires no work for data and model preparation in the first stage. In the second stage, auxiliary data $\tilde{x}^{\text{Target}}$ is used to initiate query requests to the target party to obtain the intermediate features $\tilde{H}^{\text{Target}}$ output by the target model $f_{\text{Target}}$. Subsequently, the obtained $\tilde{H}^{\text{Target}}$ and the corresponding $\tilde{x}^{\text{Target}}$ are used to train InverNet $g$, as shown in 3, where $m$ represents the number of samples in the dataset. Then, in the final stage of UIFV, the trained InverNet $g$ can be used for the reconstruction of the target data $x^{target}$. The complete attack algorithm in the Query Attack scenario is shown in Algorithm 1.
1 Function QueryAttack($f_{\text{target}}$, $H^{\text{target}}$, $D^{\text{aux}}$):
@@
3      $g_{\text{target}}$=TrainInverseNet($D^{\text{aux}}$ , $f_{\text{target}}$)
4      $\hat{x}_{\text{target}}$ = Inverse($g_{\text{target}}$, $H^{\text{target}}$)
@@
8Function TrainInverseNet($D^{\text{aux}}$, $f_{\text{target}}$):
@@
12            Randomly sample $x_{1},x_{2}...x_{k}$ from $D^{\text{aux}}$
13            Obtain $H_{i}$ by querying $f_{\text{target}}$ with $x_{i}$
14            $L(g_{\text{target}})=\frac{1}{m}\sum_{i=1}^{m}\left\|g_{\text{target}}(H_{i})-x_{i}\right\|^{2}$
15            $\theta_{g}^{(n+1)}=\theta_{g}^{(n)}-\epsilon\frac{\partial L(g_{\text{target}}^{(n)})}{\partial\theta_{g}^{(n)}}$
@@ 16            $n+=1$
18      return $g_{\text{target}}^{(\text{NIters})}$
@@
21Function Inverse($g_{\text{target}}$, $H^{\text{target}}$):
22       $\hat{x}_{\text{target}}$ = $g_{\text{target}}(H^{\text{target}})$
@@ 5.2 Data Passive Attack (DPA)
In the scenario of Data Passive Attack, the attacker’s capabilities are limited to only possessing the i.i.d. auxiliary dataset $\mathcal{D}^{\text{aux}}$ without the ability to query the target party. This could be because the target party only participates in the VFL training process but not the inference stage. In such an attack scenario, Under this assumption, the Model Preparation function in the UIFV framework needs to construct a shadow model $\tilde{f}_{\text{target}}$ that mimics the behavior of the target model.
One approach to building a shadow model is to recover the structure and parameters of the target model by querying the black-box model, allowing the shadow model to mimic the behavior of the target model Oh et al. (2019); Tramèr et al. (2016); Wang and Gong (2018). However, since we cannot query the target model, we turn to the VFL architecture and attempt to build a model that behaves similarly to the target model within the VFL framework. In this process, we require the cooperation of other participants, querying them with i.i.d. data to obtain the corresponding intermediate features. The optimization objective of the shadow model is
$\arg\min_{{\tilde{f}}}\frac{1}{m}\sum_{i=1}^{m}\mathcal{L}\left(f_{\text{top}}\left(H^{1},\cdots,\tilde{f}_{\text{target}}(x_{i}),\cdots,H^{K}\right),y_{i}\right),$ (4)
where $m$ represents the number of samples in the dataset, and $H^{k}$ is the intermediate features output of other participants, which remains constant during the training process. $\mathcal{L}$ is consistent with the loss function of the top model. This means that the training of $\tilde{f}_{\text{target}}$ is guided by the supervision of $f_{\text{top}}$.
Once the shadow model $\tilde{f}_{\text{target}}$ is trained in the first stage, it can be used in the role of $f_{\text{target}}$ together with the dataset $\mathcal{D}^{\text{aux}}$ to train the InverNet $g$ in the second stage and then to reconstruct the private data in the third stage of the UIFV framework, as in the Query Attack scenario. The complete attack algorithm in the Data Passive Attack scenario is shown in Algorithm 2.
1 Function DataPassiveAttack($f_{\text{target}}$, $H^{\text{target}}$, $D^{\text{aux}}$):
@@
3      $\tilde{f}_{\text{target}}$ = TrainShadowModel($D^{\text{aux}}$, $f_{\text{top}}$)
4      $g_{\text{target}}$ = TrainInverseNet($D^{\text{aux}}$ , $\tilde{f}_{\text{target}}$)
5      $\hat{x}_{\text{target}}$ = Inverse($g_{\text{target}}$, $H^{\text{target}}$)
@@
9Function TrainShadowModel($D^{\text{aux}}$ , $f_{\text{target}}$):
@@
13            Randomly sample $x_{1}$, $x_{2}$, … ,$x_{k}$ and labels $y_{1}$, $y_{2}$, … ,$y_{k}$ from $D^{\text{aux}}$
14            $\hat{y}_{i}$ = $f_{\text{top}}\left(H^{1},\cdots,\tilde{f}_{\text{target}}(x_{i}),\cdots,H^{K}\right)$
15            $L(\tilde{f}_{\text{target}})=\tfrac{1}{k}\sum_{i=1}^{k}y_{i}\hat{y}_{i}+(1-y_{i})(1-\hat{y}_{i})$
16            $\theta_{\tilde{f}}^{(n+1)}=\theta_{\tilde{f}}^{(n)}-\epsilon\frac{\partial L(\tilde{f}_{\text{target}}^{(n)})}{\partial\theta_{\tilde{f}}^{(n)}}$
@@ 17            $n+=1$
19      return $\tilde{f}_{\text{target}}^{({\text{NIters})}}$
@@ 5.3 Isolated Query Attack (IQA)
The most straightforward approach to creating fake data is random generation; for example, sampling pure noise from a standard Gaussian distribution as in He et al. (2019) for image data. However, tabular data, which is significantly different in distribution from image data, consists of heterogeneous features and lacks spatial or semantic relationships, making it more complex to discover and utilize relationships Borisov et al. (2022). Therefore, using randomly generated data for query requests may lead to poor reconstruction results, as verified in Section 6.4.2.
To address this issue, we introduce prior knowledge to guide the process of generating random data, aiming to enhance the quality of data reconstruction. For image data, we focus on generating smoother random samples by reducing the influence of noise and outliers in random data through sampling from a standard Gaussian distribution. For tabular data, based on the analysis in Borisov et al. (2022), we utilize the header information of the target dataset to construct pseudo data, making the fabricated data closer to the distribution of real data. For categorical variables, we adopt a one-hot encoding approach, randomly selecting a category to set its corresponding column to 1, while keeping other columns at 0. For continuous variables, we first estimate the range of values based on experience and then perform random sampling within the estimated range using a uniform distribution. We call this process the data generation module.
1 Function IsolatedQueryAttack($f_{\text{target}}$, $H^{\text{target}}$, $T$):
@@ 3      $D^{\text{fake}}$ = DataGeneration($T$)
4      $g_{\text{target}}$ = TrainInverseNet($D^{\text{fake}}$ , $f_{\text{target}}$)
5      $\hat{x}_{\text{target}}$ = Inverse($g_{\text{target}}$, $H^{\text{target}}$)
@@ 5.4 Stealth Attack (SA)
In the previous three attack scenarios, the attacker can possess an i.i.d. auxiliary dataset and/or make queries to obtain the intermediate features but does not know about the target party’s private information. In this scenario, even if the attacker loses both the auxiliary data and the query capabilities, data reconstruction is still possible if the attacker $\mathbb{P}_{\text{adv}}$ may acquire a minimal amount of the target party’s private data and explicitly knows that these data have already been used in the VFL process. For instance, $\mathbb{P}_{\text{adv}}$ might collude with some internal employees of the target party (whose data are jointly maintained by both parties) to secretly acquire a small subset of data, which is considered possible as has been noted in some studies on analogous VFL scenarios, such as Zeng et al. (2023). The knowledge of some private data used by the target party allows the attacker to start from the second stage of the UIFV framework to directly train InverNet $g$ with the objective function simplified to 5:
$\arg\min_{\theta_{g}}\left\|g(H^{k}))-x^{k}\right\|^{2}$ (5)
In this equation, $x^{k}$ signifies the secretly acquired data, while $H^{k}$ represents the corresponding intermediate features of these secret data. Then, the trained $g$ will be used in the third stage to complete private data reconstruction. The complete attack algorithm in the Stealth Attack scenario is shown in Algorithm 4.
1 Function StealthAttack($H^{\text{target}}$, $D^{\text{priv}}$):

### v2

@@ 4 UIFV Framework Overview
We have developed a comprehensive framework, the Unified InverseNet Framework in VFL (UIFV). This framework enables attackers with varying capabilities, acting as the active party, to train an InverseNet and leverage the intermediate features of the passive party during normal VFL training or inference processes to reconstruct the private data of the passive participant.
@@
In the VFL framework, when we consider the $k$-th participant ($k\neq 1$) as our attack target, we treat the bottom model $f_{k}$ of participant $\mathbb{P}_{k}$ as a feature extractor. The bottom model $f_{k}$ generates the intermediate features $H^{k}$ that are fed into the top model $f_{\text{top}}$. Considering that the intermediate layers of neural networks retain rich semantic information of input data, the proposed UIFV framework focuses on training an InverNet $g$ that establishes the relationship $g(f_{k}(x^{k}))=x^{k}$, which can then be used to upsample and reconstruct the private data $x^{k}$. The objective function for training the InverNet $g$ can be formulated as:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\arg\min_{\theta_{g}}\left\|g(H^{k})-x^{k}\right\|^{2},}$ (6)
which indicates that the private data $x^{k}$ and the bottom model $f_{k}$ are required for training the InverNet. However, such information cannot be directly obtained in realistic VFL scenarios. Therefore, Data Preparation and Model Preparation are two key functional modules in the proposed UIFV framework that respectively prepare the model and data information needed for training the InverNet. Based on whether these two modules are utilized, the attack scenarios are categorized into four types, which will be discussed in sections 5.
@@
Similarly, different methods can be employed by the Model Preparation module based on the attacker’s capabilities. If the attacker can make inference queries to the target party, which is reasonable in the VFL architecture, then the model output $f_{k}(x^{k})$ can be obtained from the queries. If the attacker cannot make query requests to the bottom model, then the attacker may train a shadow model $\tilde{f}_{k}$ to replace the model $f_{k}$ for training the InverNet. Note that in the following sections, we will omit the index $k$ of the target party when there is no ambiguity.
@@
As depicted in Fig. 4, the UIFV framework encompasses three distinct stages. The Data Preparation and Model Preparation functions are performed in the initial stage to prepare the pertinent data and model needed for training the InverNet, which is then conducted in the second stage. Then, in the final stage, the pre-trained InverNet is employed to reconstruct the private data used for model training by the target party. For different types of data, we have designed different InverNet architectures, as shown in Figure 5.
Figure 5: InverNet Architecture for Different Data Types.
@@ 5.1 Query Attack (QA)
In the Query Attack scenario, the attacker has no access to the target party’s model parameters and model gradient information during the training process, which is consistent with the black-box attack setting [7]. On the other hand, the attacker can freely initiate query requests to the target party and possesses the auxiliary dataset $D^{\text{aux}}$ , that is, i.i.d. with the target party’s private data.
1 Function QueryAttack(${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{k}}}$, $H^{\text{target}}$, $D^{\text{aux}}$):
@@
3      $g$=TrainInverNet($D^{\text{aux}}$ , ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{k}}}$)
4      $\hat{x}_{\text{target}}$ = Inverse($g$, $H^{\text{target}}$)
@@
8Function TrainInverNet($D^{\text{aux}}$, ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{k}}}$):
@@
12            Randomly sample $\tilde{x}_{1},\tilde{x}_{2}...\tilde{x}_{m}$ from $D^{\text{aux}}$
13            Obtain $\tilde{x}_{i}$ by querying ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{k}}}$ with $x_{i}$
14            $L(g)=\frac{1}{m}\sum_{i=1}^{m}\left\|g(H_{i})-x_{i}\right\|^{2}$
15            $\theta_{g}^{(n+1)}=\theta_{g}^{(n)}-\epsilon\frac{\partial L(g^{(n)})}{\partial\theta_{g}^{(n)}}$
@@ 16            $n+=1$
18      return $g^{(\text{NIters})}$
@@
21Function Inverse($g$, $H^{\text{target}}$):
22       $\hat{x}_{\text{target}}$ = $g(H^{\text{target}})$
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}g=\arg\min_{\theta_{g}}\frac{1}{m}\sum_{i=1}^{m}\left\|g(\tilde{H}_{i})-\tilde{x}_{i}\right\|^{2}}$ (7)
In this attack scenario, since the attacker owns an auxiliary dataset and is able to query the target model, the UIFV framework requires no work for data and model preparation in the first stage. In the second stage, auxiliary data $\tilde{x}$ is used to initiate query requests to the target party to obtain the intermediate features $\tilde{H}$ output by the target model $f_{k}$. Subsequently, the obtained $\tilde{H}$ and the corresponding $\tilde{x}$ are used to train InverNet $g$, as shown in 7, where $m$ represents the number of samples in the dataset. Then, in the final stage of UIFV, the trained InverNet $g$ can be used for the reconstruction of the target data $x^{target}$. The complete attack algorithm is detailed in Algorithm 1.
@@ 5.2 Data Passive Attack (DPA)
In the scenario of Data Passive Attack, the attacker’s capabilities are limited to only possessing the i.i.d. auxiliary dataset $\mathcal{D}^{\text{aux}}$ without the ability to query the target party. This could be because the target party only participates in the VFL training process but not the inference stage. In such an attack scenario, Under this assumption, the Model Preparation function in the UIFV framework needs to construct a shadow model $\tilde{f}_{k}$ that mimics the behavior of the target model.
One approach to building a shadow model is to recover the structure and parameters of the target model by querying the black-box model, allowing the shadow model to mimic the behavior of the target model [28, 29, 30]. However, since we cannot query the target model, we turn to the VFL architecture and attempt to build a model that behaves similarly to the target model within the VFL framework. In this process, we require the cooperation of other participants, querying them with i.i.d. data to obtain the corresponding intermediate features. The optimization objective of the shadow model is
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\arg\min_{{\tilde{f}_{k}}}\frac{1}{m}\sum_{i=1}^{m}\mathcal{L}\left(f_{\text{top}}\left(H^{1},\cdots,\tilde{f}_{k}(x_{i}),\cdots,H^{K}\right),y_{i}\right),}$ (8)
where $m$ represents the number of samples in the dataset, and $H^{i}(i\neq k)$ is the intermediate features output of other participants, which remains constant during the training process. $\mathcal{L}$ is consistent with the loss function of the top model. This means that the training of $\tilde{f}_{k}$ is guided by the supervision of $f_{\text{top}}$.
1 Function DataPassiveAttack(${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{k}}$, $H^{\text{target}}$, $D^{\text{aux}}$):
@@
3      $\tilde{f}_{k}$ = TrainShadowModel($D^{\text{aux}}$, $f_{\text{top}}$)
4      $g$ = TrainInverNet($D^{\text{aux}}$ , $\tilde{f}_{k}$)
5      $\hat{x}_{\text{target}}$ = Inverse($g$, $H^{\text{target}}$)
@@
9Function TrainShadowModel($D^{\text{aux}}$ , ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{k}}$):
@@
13            Randomly sample $\tilde{x}_{1}$, $\tilde{x}_{2}$, … ,$\tilde{x}_{m}$ and labels $\tilde{y}_{1}$, $\tilde{y}_{2}$, … ,$\tilde{y}_{m}$ from $D^{\text{aux}}$
14            $\hat{y}_{i}$ = $f_{\text{top}}\left(H^{1},\cdots,\tilde{f}_{k}(\tilde{x}_{i}),\cdots,H^{K}\right)$
15            $L(\tilde{f}_{k})=\tfrac{1}{m}\sum_{i=1}^{m}\tilde{y}_{i}\hat{y}_{i}+(1-\tilde{y}_{i})(1-\hat{y}_{i})$
16            $\theta_{\tilde{f}_{k}}^{(n+1)}=\theta_{\tilde{f}_{k}}^{(n)}-\epsilon\frac{\partial L(\tilde{f}_{k}^{(n)})}{\partial\theta_{\tilde{f}_{k}}^{(n)}}$
@@ 17            $n+=1$
19      return $\tilde{f}_{k}^{({\text{NIters})}}$
@@
Once the shadow model $\tilde{f}_{k}$ is trained in the first stage, it can be used in the role of ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{k}}$ together with the dataset $\mathcal{D}^{\text{aux}}$ to train the InverNet $g$ in the second stage and then to reconstruct the private data in the third stage of the UIFV framework, as in the Query Attack scenario. The complete attack algorithm in the Data Passive Attack scenario is shown in Algorithm 2.
@@ 5.3 Isolated Query Attack (IQA)
1 Function IsolatedQueryAttack(${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{k}}$, $H^{\text{target}}$, $T$):
@@ 3      $D^{\text{fake}}$ = DataGeneration($T$)
4      $g$ = TrainInverNet($D^{\text{fake}}$ , ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{k}}$)
5      $\hat{x}_{\text{target}}$ = Inverse($g$, $H^{\text{target}}$)
@@
The most straightforward approach to creating fake data is random generation; for example, sampling pure noise from a standard Gaussian distribution as in [26] for image data. However, tabular data, which is significantly different in distribution from image data, consists of heterogeneous features and lacks spatial or semantic relationships, making it more complex to discover and utilize relationships [31]. Therefore, using randomly generated data for query requests may lead to poor reconstruction results, as verified in Section 6.5.2.
To address this issue, we introduce prior knowledge to guide the process of generating random data, aiming to enhance the quality of data reconstruction. For image data, we focus on generating smoother random samples by reducing the influence of noise and outliers in random data through sampling from a standard Gaussian distribution. For tabular data, based on the analysis in [31], we utilize the header information of the target dataset to construct pseudo data, making the fabricated data closer to the distribution of real data. For categorical variables, we adopt a one-hot encoding approach, randomly selecting a category to set its corresponding column to 1, while keeping other columns at 0. For continuous variables, we first estimate the range of values based on experience and then perform random sampling within the estimated range using a uniform distribution. We call this process the data generation module.
@@ 5.4 Stealth Attack (SA)
1 Function StealthAttack($H^{\text{target}}$, $D^{\text{leak}}$):